OpenStudy (anonymous):

The value of x in the proportion 1/2:2/3 = 3/4:x is A)1 B)4/9 C)1 7/9 D)1/4

OpenStudy (anonymous):

\[\frac{ .5 }{ .75 }=\frac{ .666 }{ x }\]\[.5\times x=.5x\]\[.75\times.666=.5\]\[.5x=.5\]\[.5x \div.5=x\]\[.5\div.5=1\]\[x=1\]

OpenStudy (anonymous):

@QuestionsQuestions, I answered you question.

OpenStudy (anonymous):

Ask me if there is anything you don't understand.

OpenStudy (anonymous):

no that's perfect thank you

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

Do you need help with anything else?

OpenStudy (anonymous):

I'll see hold on

OpenStudy (anonymous):

If a/b=p/q, then a + b/b = p + b/q. True False

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

All of the following are equivalent to 3:8=x:32 except 8:3 = 32:x 3:x = 8:32 x:8 = 3:32 8:32 = 3:x

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (anonymous):

\[\frac{ 3 }{ x }=\frac{ 8 }{ 32 }\]-------- A.\[\frac{ 8 }{ 32 }=\frac{ 3 }{ x }\]-------- B.\[\frac{ 3 }{ 8 }=\frac{ x }{ 32 }\]--------- C.\[\frac{ x }{ 3 }=\frac{ 8 }{ 32 }\]---------- D.\[\frac{ 8 }{ 3 }=\frac{ 32 }{ x }\]

OpenStudy (anonymous):

I'll solve it in a minute, but I need them in proportion form for me to solve it easier.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

All others, x=12. In C, x=.75

OpenStudy (anonymous):

ooooooohhhhhhhh ok